OpenStudy (anonymous):

cos(3x)=-1 Part I: Solve this equation for 3x, finding all values on the interval [0,2π] that satisfy the equation. Part II: Use algebra to find all values of x between 0 and 2π that satisfy this equation.

OpenStudy (tkhunny):

Where does the cosine take on the value -1?

OpenStudy (anonymous):

pi?

OpenStudy (tkhunny):

You are nearly done with this problem.

OpenStudy (anonymous):

Hmm, so how would I progress from there?

OpenStudy (solomonzelman):

nvm, it can be -1 and 1, but not more

OpenStudy (tkhunny):

You have a solution, 3x = pi. You know the cosine is periodic with period 2pi \(3x = \pi + 2k\pi\) for any integer \(k\). Divide by 3. Figure out what 'k' means.

OpenStudy (anonymous):

Hmm, I just can't figure out what 'k' could mean, could you give me a bit of a hint? x)

OpenStudy (tkhunny):

Did already. It is any integer. Your task is to find which integers put the solution in \([0,2\pi]\).

OpenStudy (anonymous):

Hm would it be something like -1?

OpenStudy (tkhunny):

?? You have the equation already: \(3x = \pi + 2k\pi\) Solve for x. \(x = \dfrac{\pi}{3} + \dfrac{2k\pi}{3} = \dfrac{\pi}{3}\left(1 + 2k\right)\) Solve for k \(1+2k = 0\) \(1+2k = 1\) \(1+2k = 2\) \(1+2k = 3\) \(1+2k = 4\) \(1+2k = 5\) \(1+2k = 6\) Why did I quit at 6?